# A Dimca-Greuel type inequality for foliations

### v2

@@ 3. Reduction of singularities of foliations and their separatrices
(16) $m_{q^{(j)}}(\mathcal{F}^{(j)})=\left\{\begin{array}[]{ll}\nu_{q^{(j)}}(\mathcal{F}^{(j)})+1&\hbox{\rm if $\sigma_{j+1}$ is dicritical}\\ \nu_{q^{(j)}}(\mathcal{F}^{(j)})&\hbox{\rm if $\sigma_{j+1}$ is non-dicritical}.\\ \end{array}\right.$
Notice that for $j=0$ we get $q^{(0)}=p$, $\mathcal{F}^{(0)}=\mathcal{F}$ and the equality (16) for $j=0$ becomes:
(17) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical},\\ \end{array}\right.$
where $\sigma$ is a blow-up at the point $p$.
The following lemma relates $R_{p}(\mathcal{F},B)$ to the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ (respectively $\mathcal{F}^{(N)}$) with respect to $B$ (respectively $B^{(N)}$).
Lemma 3.1.
With the above notations we get
$R_{p}(\mathcal{F},B)=GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right].$
Proof.
From blow-up formula of the GSV index [4, p.30] we get
(18) $GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})$ $=$ $GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})$
$+$ $\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})),\,\,\,\,\hbox{\rm for }\,\,\,0\leq j\leq N-1.$
Up substituting the equality (18) in (15), we get
$R_{p}(\mathcal{F},B)$ $=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right]$
$=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}))\right]$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
$=$ $\sum_{j=0}^{N-1}\left[GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})-GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})\right]$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
$=$ $GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right],$
and the lemma follows. ∎
For a germ of a holomorphic foliation $\mathcal{F}$ and a separatrix $B\in\text{Sep}_{p}(\mathcal{F})$ at $(\mathbb{C}^{2},p)$, we introduce the number $\text{DG}_{p}(\mathcal{F},B)$ as follows:
(21) $\text{DG}_{p}(\mathcal{F},B)$ ::=:= $\left\{\begin{array}[]{ll}R_{p}(\mathcal{F},B)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})&\hbox{\rm if $B\in\text{ISO}_{p}(\mathcal{F})$}\\ R_{p}(\mathcal{F},B)&\hbox{\rm if $B\in\text{Dic}_{p}(\mathcal{F})$}.\\ \end{array}\right.$
From (14), we deduce that $\alpha_{i}$ are analytic invariants, thus we have $\text{DG}_{p}(\mathcal{F},B)$ is also an analytic invariant depending of $\mathcal{F}$ and $B$. As a consequence of Lemma 3.1, we can establish $\text{DG}_{p}(\mathcal{F},B)$ in terms of the GSV-indices of the foliations $\mathcal{F}$ and $\mathcal{F}^{(N)}$ and some numbers depending of the resolution of the singularities of $B$.
4. Proof of Theorem A
First of all, we will study the behaviour of $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ under a blow-up $\sigma$ at $p\in\mathbb{C}^{2}$. For this, we consider $\alpha_{0}$ as in (14) for $j=0$.
We have the following lemma.
Lemma 4.1.
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B\in\text{Sep}_{p}(\mathcal{F})$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) by a blow-up $\sigma$ at $p$ then
$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0},$
where $q\in\tilde{B}\cap\sigma^{-1}(p)$, and $\alpha_{0}$ is as in (14).
Proof.
By [6, Equality (1.2), p. 291] we get
where $m_{p}(\mathcal{F})$ was introduced in (17).
On the other hand by the behaviour under blow-up of the GSV index (see [4, p.30]) we get
@@ 4. Proof of Theorem A
and after [8, Proposition 6.2] we have
@@
From [19, Equation (4)] we obtain
(25) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha_{0},$
where $\alpha_{0}$ was defined in (14). From (24), (23), and (25) we deduce
(26) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha_{0}.$
By (22), (23) and (26) it holds
$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0}.$
@@
If $\mathcal{F}$ is a reduced foliation at $p\in\mathbb{C}^{2}$, and $B$ is a smooth separatrix through $p$, then, by Remark 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F}.B)=\tau_{p}(\mathcal{F},B)$. Therefore, according to [3, p. 538], we have $\tau_{p}(\mathcal{F},B)=1$, if $p$ is a non-degenerate singularity, or if $p$ is a saddle-node singularity and $B$ is the strong separatrix; and $\tau_{p}(\mathcal{F},B)=k+1$, if $p$ is a saddle-node singularity and $B$ is the weak separatrix. Here $k\in\mathbb{Z}^{+}$ is a formal invariant in the case that $p$ is a saddle-node singularity of $\mathcal{F}$. In any cases, $\tau_{p}(\mathcal{F},B)>0$.
@@ 4.1. Proof of Theorem A
Let $\mathcal{F}$ be a holomorphic foliation on $(\mathbb{C}^{2},p)$, and let $B$ be a separatrix of $\mathcal{F}$ at $p$. Then, on a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we marked the reduction process of the pair $(\mathcal{F},B)$:
(27) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ coincide with the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$. Put $\mu^{(i)}=\mu_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$, for any $i\in\{0,\ldots,N\}$.
Applying Lemma 4.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\overset{\sigma_{i+1}}{\longrightarrow}(\mathcal{F}^{(i)},B^{(i)})$, we get
$3\mu^{(i)}-4\tau^{(i)}=3\mu^{(i+1)}-4\tau^{(i+1)}+\left[\nu_{q^{(i)}}(B^{(i)})(2\nu_{q^{(i)}}(B^{(i)})-m_{q^{(i)}}(\mathcal{F}^{(i)})-1)-4\alpha_{i}\right]$
@@
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ $=$ $3\mu^{(N)}-4\tau^{(N)}$
$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
$=$ $3\mu^{(N)}-4\tau^{(N)}+R_{p}(\mathcal{F},B),$
where the last equality follows from (15).
We have two possibilities: either $B\in\text{ISO}_{p}(\mathcal{F},B)$ or $B\in\text{Dic}_{p}(\mathcal{F})$. If $B\in\text{ISO}_{p}(\mathcal{F},B)$, there exists a component $D\subset\mathcal{D}$ with $D$ non-dicritical such that $\mathcal{F}^{(N)}$ has a reduced singularity at $q^{(N)}\in\mathcal{D}\cap B^{(N)}$. In this case, we have $\mu^{(N)}=\tau^{(N)}$ by Remark 4.2, and thus
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=-\tau^{(N)}+R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
If $B\in\text{Dic}_{p}(\mathcal{F})$, then there exists a component $D\subset\mathcal{D}$ with $D$ dicritical such that $B^{(N)}$ is transversal to $D$ at $q^{(N)}$, and $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$. In this case, we get
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
In both cases, the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ implies $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows. $\Box$
The following example illustrates that the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ in Theorem A cannot be removed.
@@
where $m\in\mathbb{Z}_{>0}$. Let $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$. We have that $B$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$ such that
@@
A straightforward calculation reveals that $\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}<\frac{4}{3}$ for $60<m<$ 6, and
$\lim_{m\to+\infty}\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}=2.$
On the other hand, we compute $\text{DG}_{0}(\mathcal{F}_{m},B)$ for $m\geq 1$. Note that $\nu_{0}(\mathcal{F}_{m})=m+2$, and $\nu_{0}(B)=2m$, and using [19, equation 4] we get
$\alpha_{0}=\left\{\begin{array}[]{cl}1&\text{ if $m=1$}\\ m^{2}+m&\text{ if $m>1$}\end{array}\right.$
After a blow-up at $0\in\mathbb{C}^{2}$, the strict transform of $B$ is a smooth curve, and $m_{0}(\mathcal{F}_{m})=m+2$. The reduction of singularities of the pair $(\mathcal{F}_{m},B)$ is achieved in $2m$ steps, and $B\in\text{Dic}_{0}(\mathcal{F}_{m})$. Thus
$\text{DG}_{0}(\mathcal{F}_{1},B)$ $=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F})-1)-4\alpha_{0}\right]$
$+\left[\nu_{q^{(1)}}(B^{(1)})(2\nu_{q^{(1)}}(B^{(1)})-m_{q^{(1)}}(\mathcal{F}^{(1)})-1)-4\alpha_{1}\right]$
$=$ $\left[2(4-3-1)-4(1)\right]+\left[1(2-2-1)-4(0)\right]$
$=$ $-5$
and for $m>1$, we have
$\text{DG}_{0}(\mathcal{F}_{m},B)$ $=$ $\sum_{j=0}^{2m-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}_{m})-1)-4\alpha_{j}\right]$
$=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F}_{2m})-1)-4\alpha_{0}\right]$
$+\left[\nu_{q^{(2m-1)}}(B^{(2m-1)})(2\nu_{q^{(2m-1)}}(B^{(2m-1)})-m_{q^{(2m-1)}}(\mathcal{F}^{(2m-1)}_{2m})-1)-4\alpha_{2m-1}\right]$
$=$ $\left[2m\left(2(2m)-(m+2)-1\right)-4(m^{2}+m)\right]+\left[1(2-2-1)-4(0)\right]$
$=$ $1\displaystyle 2m^{2}-10m-$ 1 .
Therefore, $\text{DG}_{0}(\mathcal{F}_{m},B)\geq 0$ if and only if $m\geq 6$. This shows that the hypothesis of Theorem A is optimal.

### v3

@@ 3. Proof of Theorem A
$m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is nondicritical}.\\ \end{array}\right.$
On the other hand by the behavior under blow-up of the GSV index [3, p. 30] we get
@@
and after [7, Proposition 6.2] we have
@@
From [17, Equation (4)] we obtain
(15) $\tau_{p}(B)-\tau_{q}(\tilde{B})\geq\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
where $\alpha$ was defined in (10). From (14), (13), and (15) we deduce
(16) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})\leq\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha.$
After (12), (13) and (16) we get
(17) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})$$\geq$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)$
$+$ $4\alpha-\nu_{p}(B)(\nu_{p}(B)-1)$
$>$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B),$
where the last inequality follows by (10), since $\nu_{p}(B)\geq 2.$
Suppose now that $B$ is a non-singular curve. By a change of coordinates, if necessary, we can suppose that $B$ is given by $x=0$. Since $B$ is a separatrix of $\mathcal{F}:\omega=Pdx+Qdy$ then $\omega\wedge dx=Qdx\wedge dy$. Hence $Q=xh$ for some convergent power series $h(x,y)\in\mathbb{C}\{x,y\}$ such that $h(0,y)\not=0$ and $\mathcal{F}:\omega=P(x,y)dx+xh(x,y)dy$. Since $\gamma(t)=(0,t)$ is a parametrization of $B$, then $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}P(0,t)$ and
(18) $\tau_{p}(\mathcal{F},B)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,xh,x)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,x)=\mu_{p}(\mathcal{F},B).$
Moreover by (14) we have the equality $GSV_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$. Thus, if $\nu_{p}(B)=1$ then $\nu_{p}(\tilde{B})=1$, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)$ and $\mu_{p}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\tilde{\mathcal{F}},\tilde{B})$, so (11) becomes an equality. Moreover, if (11) is an equality, then $B$ is necessarily smooth. Indeed, if $B$ were singular, by (10) the number $-\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha$ is positive and from the equality (17) we have that (11) is a strict inequality that is a contradiction. Hence the lemma follows. ∎
@@
By definition, for any branch $B$ we have $\tau_{p}(B)\leq\mu_{p}(B)$, so after [7, Proposition 6.2] and [7, Propositon 5.7] we get $\tau_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)\leq\mu_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)$ and
(19) $\tau_{p}(\mathcal{F},B)\leq\mu(\mathcal{F},B).$
In addition, by (18), if $B$ is smooth then $\tau_{p}(\mathcal{F},B)=\mu(\mathcal{F},B)$.
@@ 3.1. Proof of Theorem A
Suppose that $\mathcal{F}$ is a holomorphic foliation at $p\in\mathbb{C}^{2}$, $B$ is a separatrix of $\mathcal{F}$ at $p$. First, assume that $B$ is singular, that is, $\nu_{p}(B)\geq 2$. Then, there is a sequence of blow-ups of $\mathcal{F}$ at $p$ (cf. Seidenberg [16]):
(20) $(\mathcal{F}^{(N)},B^{(N)})\longrightarrow\cdots\longrightarrow(\mathcal{F}^{(2)},B^{(2)})\longrightarrow(\mathcal{F}^{(1)},B^{(1)})\longrightarrow(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
where $B^{(i+1)}$ denotes the strict transform of $B^{(i)}$ thought $q^{(i)}$ for each $i=0,\ldots,N-1$, $B^{(N)}$ is smooth, and $\mathcal{F}^{(N)}$ is a foliation with a simple singularity at $q^{(N)}\in B^{(N)}$, or $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$, and $B^{(N)}$ is a dicritical separatrix (cf. [11, p. 1423]). Put $\mu^{(i)}:=\mu_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ for any $i\in\{1,\ldots,N\}$.
Applying Lemma 3.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\longrightarrow(\mathcal{F}^{(i)},B^{(i)})$, we get
$3\mu^{(i)}-4\tau^{(i)}+GSV_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})\leq 3\mu^{(i+1)}-4\tau^{(i+1)}+GSV_{q^{(i+1)}}(\mathcal{F}^{(i+1)},B^{(i+1)}).$
@@
$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)<3\mu^{(N)}-4\tau^{(N)}+GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})=0,$
where the last equality holds, since $B^{(N)}$ is a smooth curve.
If $B$ is a smooth curve then, by Remarks 3.2, and 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$, then the inequality (4) is an equality. Reciprocally, if (4) is an equality, then $B$ is necessarily smooth. Indeed, if $B$ were singular, by Lemma 3.1, $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)<$ 0, this is a contradiction. Hence, the theorem follows. $\Box$
4. Examples
In this section, we provide two examples of holomorphic foliations and separatrices that illustrate Theorem A.
Example 4.1.
Let $\mathcal{G}_{m,n}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
$\eta_{m,n}=mxdy-nydx,$
where $1\leq m\leq n$ are integers. The curve $B:y^{m}-x^{n}=0$ is a separatrix of $\mathcal{G}_{m,n}$ at $0\in\mathbb{C}^{2}$ and
$GSV_{0}(\mathcal{G}_{m,n},B)$ $=$ $m+n-mn,$
$\mu_{0}(\mathcal{G}_{m,n},B)$ $=$ $1,$
$\tau_{0}(\mathcal{G}_{m,n},B)$ $=$ $1$
We get
$3\mu_{0}(\mathcal{G}_{m,n},B)-4\tau_{0}(\mathcal{G}_{m,n},B)+GSV_{0}(\mathcal{G}_{m,n},B)=m+n-mn-1=-(m-1)(n-1)\leq 0$
Thus, Theorem A is verified.
@@ 4. Examples
where $m\in\mathbb{Z}_{>0}$. The curve $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We get
@@
A straightforward calculation reveals that
$3\mu_{0}(\mathcal{F}_{m},B)-4\tau_{0}(\mathcal{F}_{m},B)+GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}-2&\text{if $m=1$}\\ -6m&\text{ if $m>1$}.\end{array}\right.$
Therefore, Theorem A is verified.